A 72.0 kg box slides down a 30.0◦
ramp with
an acceleration of 3.10 m/s2
.
72 kg
µk
3.1 m/s
2
30◦
Find µk between the box and the ramp.
The acceleration of gravity is 9.81 m/s
2
.
1
Expert's answer
2019-12-16T11:04:51-0500

The acceleration will be


"a=g\sin{30}-\mu g \cos{30}"

"3.1=9.81(\sin{30}-\mu \cos{30})"

"\mu=0.212"